# PASTA: A Scalable Framework for Multi-Policy AI Compliance Evaluation

### v1

@@ 8. Discussion
In this section, we position our work in a broader perspective and discuss the wider applications and implications of PASTA across 3 domains: (1) strengthening documentation and harmonization practices, (2) supporting real-world compliance workflows and navigating regulatory complexity, and (3) enabling safe, collaborative, and educational use of AI-assisted compliance tools.
@@ 8.2. Supporting Real-World Workflows and Navigating Regulatory Complexity
Our usability study unexpectedly revealed that PASTA integrates naturally into early–mid stages of AI system development workflow. Several participants in our usablity study described it as a practical mechanism for identifying compliance obligations and missing safeguards while still exploring feasibility or refining a partially built system. Unlike prompt-based risk-prediction tools that operate at a high level, and unlike code-dependent audit tools that assume a finished product, PASTA introduces an AI compliance framework that occupies a middle ground—specific enough to surface actionable issues yet flexible enough to support iterative development or early-stage experimentation.
PASTA’s multi-dimensional design can support small, multi-role teams by providing views that align with needs of different roles. The model card functions as a authoring artifact for audits, while the high-level summaries and heatmap offer an accessible entry point for developers and designers without legal expertise. The heatmap and issue–fix tables also help product managers and stakeholders identify priorities and translate findings into concrete tasks. Participants further noted that PASTA’s structured outputs can be integrated into ticketing or audit systems, enabling distributed teams to coordinate around standardized evidence to better supporting larger team settings.
The system also serves as a software-architectural model for compliance tools seeking to adapt to future regulatory change, which is fairly meaningful in today’s fast-evolving policy landscape. As both the preprocessing pipeline and evaluation logic treat policies as modular inputs, newly introduced or amended regulations can be incorporated without redesigning the framework, corresponding to our first design goal regarding scalability. This architecture naturally supports governance tasks such as policy versioning and longitudinal compliance tracking. The ability to “plug in” future or customized policies is essential for AI systems operating across multiple jurisdictions. While commercial solutions often adopt similar modular designs, we openly share our implementation details, making this policy adaptability available for community-driven improvements.
@@ 8.3. Safe and Educational Uses of PASTA
PASTA also surfaced unanticipated value as both an educational tool. One participant, a university instructor, noted that the system could serve as a clear and approachable introduction to AI policy and RAI in academic or professional training contexts. They emphasized that the tool’s interactive nature creates a realistic and engaging way for learners to explore policies and AI governance. Students can input their own AI systems, which increases motivation as feedback directly concerns their work. The hands-on process can also help them understand policies in depth on a practical level. By adjusting the model card content, students can experiment with policy thresholds, seeing what counts as compliant, what constitutes a violation, and whether missing information is flagged or ignored. They can also examine how different policies respond to the same system, learning the specifics of policies and how they differ. Additionally, by varying the level of detail in their descriptions, learners gain a clearer understanding of the depth and clarity expected in AI compliance audits, reinforcing broader principles emphasized in AI governance. This makes PASTA potentially useful for teaching computer science students, AI developers, and even corporate employees, as organizations could upload their own internal policies for training purposes. These emergent uses highlight opportunities to extend PASTA into instructional settings, expanding its role beyond compliance evaluation alone.
@@ A.1. Model Selection
Claude was selected for PASTA evaluation process based on three key strengths: (1) superior long-context handling, with models supporting context windows of up to 200K tokens, which is critical for processing large and fragmented legal or policy documents and aligning them with model card sections  [Anthropic, 2023a]; (2) consistent and schema-aligned output generation, as Claude performs well in producing structured compliance score tables and section-by-section explanations  [Anthropic, 2023b]; and (3) seamless integration, as Claude is natively supported within LangChain, enabling efficient coordination across the multi-stage evaluation workflow.
@@ A.2. PASTA Model Card Section Adaptations
  • General Information: Expanded to include detailed versioning practices and compliance-aware change tracking according to the documentation requirements in the EU AI Act [European Parliament and Council of the European Union, 2024].
  • Intended Uses: Fully adopted from Mitchell et al.’s design of the Intended Uses section [Mitchell et al., 2019].
@@
  • Training Data Information: Combines the Evaluation Data and Training Data sections from Mitchell et al. and the Kaggle Model Card [Mitchell et al., 2019, Shankar, 2023]. This unified structure is better suited for AI systems beyond machine learning models, where the distinction between training and evaluation datasets is often less meaningful. By consolidating these sections, we emphasize a system-level perspective on data provenance, collection practices, and bias mitigation, aligning more closely with the needs of compliance evaluation and regulatory transparency.
  • System Performance and Evaluation: Expanded from the Quantitative Analyses section in the Kaggle Model Card [Shankar, 2023]. This section introduces disaggregated subgroup evaluations, detailed real-world testing contexts, adversarial testing scenarios, and robustness assessments. These additions provide a more granular view of system behavior across diverse conditions and user groups, which is critical for identifying potential compliance risks and ensuring fairness, safety, and reliability in real-world deployments.
  • Ethical Considerations: Adopted from Mitchell et al.’s design of the ethical considerations section and structured based on the description in the Kaggle Model Card [Mitchell et al., 2019, Shankar, 2023].
  • Maintenance and Monitoring: Includes detailed procedures for human oversight, real-time intervention, and systematic update tracking required by the EU AI Act [European Parliament and Council of the European Union, 2024].
@@ A.3. Sample Evaluation Result
3 { ”score”: 3, ”description”: ”The system name ’Crop Health Monitor’ may imply real-time monitoring capabilities not supported by the documented system functionality.” }
@@
5 { ”score”: 2, ”description”: ”The term ’yield forecasting’ is mentioned in the system name, but no forecasting methodology is provided, which introduces minor ambiguity regarding accuracy and compliance transparency.” }
A.4. Model Card Template
Category Section Description
A.5. Pair-wise Comparison Rubrics
Score Description
Table 12. Relevance Score Rubric (0–5) during the irrelevancy mapping stage: scores $\leq 1$ indicate negligible impact, while higher scores reflect increasing importance for compliance assessment across policy articles.
Score Description
Table 13. Violation scoring rubric (0–5) used in PASTA’s pairwise evaluation. The scale distinguishes between full compliance, ambiguous coverage, and clear violations, enabling nuanced assessment of system–policy alignment.
A.6. PASTA UI
Appendix B Technical Evaluation
The following policy articles were selected for expert evaluation and model comparison:
@@ Appendix B Technical Evaluation
Appendix C Technical Evaluation Plots

### v2

@@ 8. Discussion
In this section, we position our work in a broader perspective and discuss the wider applications and implications of PASTA across three domains: (1) strengthening documentation and harmonization practices, (2) supporting real-world compliance workflows and navigating regulatory complexity, and (3) enabling safe, collaborative, and educational use of AI-assisted compliance tools.
@@ 8.2. Supporting Real-World Workflows and Navigating Regulatory Complexity
Our usability study unexpectedly revealed that PASTA integrates naturally into early–mid stages of AI system development workflow. Several participants in our usability study described it as a practical mechanism for identifying compliance obligations and missing safeguards while still exploring feasibility or refining a partially built system. Unlike prompt-based risk-prediction tools that operate at a high level, and unlike code-dependent audit tools that assume a finished product, PASTA introduces an AI compliance framework that occupies a middle ground—specific enough to surface actionable issues yet flexible enough to support iterative development or early-stage experimentation.
PASTA’s multi-dimensional design can support small, multi-role teams by providing views that align with needs of different roles. The model card functions as an authoring artifact for audits, while the high-level summaries and heatmap offer an accessible entry point for developers and designers without legal expertise. The heatmap and issue–fix tables also help product managers and stakeholders identify priorities and translate findings into concrete tasks. Participants further noted that PASTA’s structured outputs can be integrated into ticketing or audit systems, enabling distributed teams to coordinate around standardized evidence to better support larger team settings.
The system also serves as a software-architectural model for compliance tools seeking to adapt to future regulatory change, which is fairly meaningful in today’s fast-evolving policy landscape. As both the preprocessing pipeline and evaluation logic treat policies as modular inputs, newly introduced or amended regulations can be incorporated without redesigning the framework, corresponding to our first design goal regarding scalability. This architecture naturally supports governance tasks such as policy versioning and longitudinal compliance tracking. The ability to “plug in” future or customized policies is essential for AI systems operating across multiple jurisdictions. While commercial solutions often adopt similar modular designs, we openly share our implementation details, making this policy adaptability available for community-driven improvements. We note that our technical evaluation assesses whether PASTA’s outputs are directionally aligned with expert judgment, rather than whether they meet a definitive standard of legal sufficiency. In practice, what constitutes “good enough” compliance reasoning depends on deployment context, organizational risk tolerance, and regulatory stakes, reinforcing that PASTA is designed to support decision-making rather than replace formal legal review.
@@ 8.3. Safe and Educational Uses of PASTA
PASTA also surfaced unanticipated value as an educational tool. One participant, a university instructor, noted that the system could serve as a clear and approachable introduction to AI policy and RAI in academic or professional training contexts. They emphasized that the tool’s interactive nature creates a realistic and engaging way for learners to explore policies and AI governance. Students can input their own AI systems, which increases motivation as feedback directly concerns their work. The hands-on process can also help them understand policies in depth on a practical level. By adjusting the model card content, students can experiment with policy thresholds, seeing what counts as compliant, what constitutes a violation, and whether missing information is flagged or ignored. They can also examine how different policies respond to the same system, learning the specifics of policies and how they differ. Additionally, by varying the level of detail in their descriptions, learners gain a clearer understanding of the depth and clarity expected in AI compliance audits, reinforcing broader principles emphasized in AI governance. This makes PASTA potentially useful for teaching computer science students, AI developers, and even corporate employees, as organizations could upload their own internal policies for training purposes. These emergent uses highlight opportunities to extend PASTA into instructional settings, expanding its role beyond compliance evaluation alone.
@@ 10. Conclusion
Acknowledgements.
This research was supported by the Korea Institute of Science and Technology (KIST) institutional program (26E0062). We sincerely appreciate their support, which made this work possible.
@@ A.1. Model Selection
Claude was selected for PASTA evaluation process based on three key strengths: (1) superior long-context handling, with models supporting context windows of up to 200K tokens, which is critical for processing large and fragmented legal or policy documents and aligning them with model card sections  (Anthropic, 2023b); (2) consistent and schema-aligned output generation, as Claude performs well in producing structured compliance score tables and section-by-section explanations  (Anthropic, 2023a); and (3) seamless integration, as Claude is natively supported within LangChain, enabling efficient coordination across the multi-stage evaluation workflow.
@@ A.2. PASTA Model Card Section Adaptations
  • System Data Information: Combines the Evaluation Data and Training Data sections from Mitchell et al. and the Kaggle Model Card (Mitchell et al., 2019; Shankar, 2023). This unified structure is better suited for AI systems beyond machine learning models, where the distinction between training and evaluation datasets is often less meaningful. By consolidating these sections, we emphasize a system-level perspective on data provenance, collection practices, and bias mitigation, aligning more closely with the needs of compliance evaluation and regulatory transparency.
  • System Performance and Evaluation: Expanded from the Quantitative Analyses section in the Kaggle Model Card (Shankar, 2023). This section introduces disaggregated subgroup evaluations, detailed real-world testing contexts, adversarial testing scenarios, and robustness assessments. These additions provide a more granular view of system behavior across diverse conditions and user groups, which is critical for identifying potential compliance risks and ensuring fairness, safety, and reliability in real-world deployments.
  • Ethical Considerations: Adopted from Mitchell et al.’s design of the ethical considerations section and structured based on the description in the Kaggle Model Card (Mitchell et al., 2019; Shankar, 2023).
  • Maintenance and Monitoring: Includes detailed procedures for human oversight, real-time intervention, and systematic update tracking required by the EU AI Act (European Parliament and Council of the European Union, 2024).
@@ A.3. Sample Evaluation Result
3 { ”score”: 3, ”description”: ”The system name ‘Crop Health Monitor‘ may imply real-time monitoring capabilities not supported by the documented system functionality.” }
@@
5 { ”score”: 2, ”description”: ”The term ‘yield forecasting‘ is mentioned in the system name, but no forecasting methodology is provided, which introduces minor ambiguity regarding accuracy and compliance transparency.” }
Appendix B Technical Evaluation and Pair-wise Comparison Scoring Rubrics
B.1. Curated Articles for Evaluation
The following policy articles were selected for expert evaluation:
@@ B.1. Curated Articles for Evaluation
B.2. Pair–wise Comparison Scoring Rubrics
Table 11. Relevance Score Rubric (0–5) during the irrelevancy mapping stage. Scores $\leq 1$ indicate negligible impact, while higher scores reflect increasing importance for compliance assessment across policy articles.
Score Description
Table 12. Violation scoring rubric (0–5) used in PASTA’s pairwise evaluation. The scale distinguishes between full compliance, ambiguous coverage, and clear violations.
Score Description
B.3. Technical Evaluation Plots
@@
B.4. PASTA UI
B.5. Model Card Template
Category Section Description